Proceedings of the V Brazilian Conference on Neural Networks - V Congresso Brasileiro de Redes Neurais pp. 123–124, April 2–5, 2001 - PUC, Rio de Janeiro - RJ - Brazil

L^{A}TEX Guidelines for V CBRN Proceedings Manuscripts

Isaac Newton^{1}, Albert Einstein^{2} 1Newton’s affiliation and address 2Einstein’s affiliation and address E-mails: isaac newton@trin.cam.ac.uk, albert einstein@ias.edu

Abstract

The ABSTRACT is to be in fully-justified italicized text, at the top of the left-hand column, below the author and affiliation information. Its first line should be indented 0.5 cm. Use the word “Abstract” as the title, in 12-point Times, boldface type, centered relative to the column, initially capitalized. The abstract is to be in 10point, single-spaced type, with 50-200 words. Leave two blank lines after the Abstract, then begin the main text. The Abstract and Main Text are to be in a two-column format. Even if the article is written in Portuguese or Spanish, the abstract should be written in English only, that is, do not include abtracts in other languages.

1. Introduction

1.1. Formatting your paper

The manuscript should be written on A4 paper, in two columns, with up to 6 pages. Columns are to be 8.0 cm wide with a 1.0 cm space between them. On all pages the top and bottom margins should be 2.5 cm and the left and right margins should be 2.0 cm from the edge of the page.

1.2. Page numbering and header

To ease the editoring of the electronic version of the proceedings, please include a header in the first page and number all pages.

All pages should be numbered sequentially, starting from the first page number assigned to your article. Page numbers should be centered at the bottom of the page. The distance between the bottom of the number and the edge of the page should be 1.5 cm.

Only the first page should have a header. The text of the header should be as shown in this example (use the page numbers assigned to your article), whatever the language of the article. This header should be right-justified. The distance between the top of the header and the edge of the page should be 1.5 cm.

Page numbers and header should be in 10-point Times.

1.3. Type-style and fonts

Wherever Times is specified, Times Roman or Times New Roman may also be used. If none is available on your word processor, please use the font closest in appearance to Times that you have access to.

MAIN TITLE. The title should be in 14-point Times, boldface type. Capitalize the first letter of nouns, pronouns, verbs, adjectives, and adverbs; do not capitalize articles, coordinate conjunctions, or prepositions (unless the title begins with such a word). Leave two blank lines after the title.

AUTHOR NAME(s) and AFFILIATION(s) are to be centered beneath the title and printed in 12-point Times, non-boldface type. This information is to be followed by two blank lines.

MAIN TEXT. Type main text in 10-point Times, single-spaced. Do NOT use double-spacing. All paragraphs should be indented 0.5 cm. Make sure your text is fully justified—that is, flush left and flush right. Please do not place any additional blank lines between paragraphs. The figure and table captions should be in 10point Times. Initially capitalize only the first word of first, second, and third-order headings.

FIRST ORDER HEADINGS (for example, 1. Introduction) should be 12-point Times boldface, initially capitalized, flush left, with one blank line before, and one blank line after.

SECOND ORDER HEADINGS (for example, 1.1. Database elements) should be 11-point Times boldface, initially capitalized, flush left, with one blank line before, and one after. If you require a third-order heading (we discourage it), use 10-point Times, boldface, initially capitalized, flush left, preceded by one blank line, followed by a period and your text on the same line.

Table 1: Table of prices

Price until date Price after date
Student  x  y
Professional  z  k

1.4. Footnotes

Please use footnotes sparingly^{1} and place them at the bottom of the column on the page on which they are referenced. Use 8-point Times type, single-spaced.

1Example of footnote.

Figure 1: Caption of figure

1.5. Illustrations and graphics

All illustrations and graphics should be black and white or gray-leveled, and centered. We discourage the use of coloured figures, since not everyone has access to high-quality coloured printers/xerox machines.

1.6. Equations

All equations should be centered, like eq. (1), and typed in 10-point Times. Equation numbers should be enclosed by parenthesis and right-justified.

E = mc^{2}  (1)

1.7. Bibliographical References

List and number all bibliographical references in 9point Times, single-spaced, at the end of your paper, in order of appearance. When referenced in the text, enclose the citation number in square brackets, for example [1], [2], and [3]. If necessary, include the name(s) of editors of referenced books.

1.8. L^{A}TEX specific instructions

This example was processed using a L^{A}TEX 2ε installation. If you are using L^{A}TEX 2.09, please read the instructions at the beginning of the style file 5cbrn inst.sty.

If you are writting in another language rather than English, please uncomment the corresponding line at the beginning of the L^{A}TEX template file 5cbrn inst.tex to enable the use of the babel package.

It is essential to instruct L^{A}TEX to use Type 1 fonts in the text, instead of the default Type 3 (bitmapped) fonts. This will make the text in your article scalable and more readable. An easy way to obtain Type 1 fonts is to use the Times package [4, p.340], as it is done in this example.

1.9. Submission instructions

The final version of the article should be in the PDF format. The file should be named 5cbrn xyz.pdf (xyz is the reference number for your article) and should be placed, using any ftp client program, in the following machine/directory:

ftp://labattmot.ele.ita.br/5cbrn/

Set your ftp client program to ”anonymous user” and (this is very important) ”bin transfer mode”.

The PDF file should be generated at 300-dpi resolution for A4 (210 x 297 mm) paper size. This is very important since the great majority of printers use 300-dpi resolution and 600-dpi printers are still very expensive.

One possible way to generate your article in the PDF format directly from the L^{A}TEX file is to use the PDFTeX/PDFLaTeX package (see http://www.tug.org/applications/pdftex/). Another (and perhaps easier) possible procedure is as follows: 1) generate the postscript file using the

command  dvips -t a4 -D300  file.dvi

-o file.ps, 2) use the freeware GhostView (see http://www.ele.ita.br/software/#Ghostview or http://www.cs.wisc.edu/˜ ghost/) to generate the PDF file from the Postscript file (for instance, in GsView 2.7 for MS-Windows: load the .ps file and select the options ”File/Print/Device: pdfwrite/Resolution: 300-dpi/Print to file”).

Before placing the PDF file with your article in the conference anonymous ftp server, check your PDF file using the freeware Adobe Acrobat Reader 4.0 (freely available at the Adobe web site (http://www.adobe.com/ or http://www.adobe.com.br/) to view and print your article.

2. Conclusions

For more help and hints about the manuscript format and how to produce the PDF file with your article, see the section ”Instructions for Authors” at the CNRN web site http://www.ele.ita.br/cnrn/.

Please direct any L^{A}TEX specific questions to the CBRN-LaTeX Mailing List (e-mail cbrn-latex@ele.ita.br ).

References

[1] A. S. Weigend and N. A. Gershenfeld. Results of the time series prediction competition at the Santa Fe Institute. In Proceedings of the III International Conference on Neural Networks, San Francisco, pages 1786–1793, 1993.

[2] S. Shah, F. Palmieri, and M. Datum. Optimal filtering algorithms for fast learning in feedforward neural networks. Neural Networks, 5(5):779–787, 1992.

[3] J. M. Zurada. Introduction to Artificial Neural Systems. West Publishing Company, New York, 1992.

[4] M. Goossens, F. Mittelbach, and A. Samarin. The LaTeX Companion. Addison-Wesley, Reading, MA, 1994.